OpenStudy (anonymous):

If y varies inversely as x and y = 4 when x = 9, what is x when y = 72?

OpenStudy (anonymous):

x=1/2

OpenStudy (anonymous):

x is increased by 8 times because 9*8=72, so y decreases by 8 times, y=4/8=1/2.

OpenStudy (anonymous):

that's what the word "inversely" means.